Site Conditions:
  • The site was drier than yesterday, but still muddy in places.
  • No smoke was observed, but the odor was still prevalent.
  • More dirt was added at the top of the hill on the southern area, and dirt work continued on the northern area.  
  • EPA said 188 truckloads of material were brought in from the borrow pit on Tuesday, and they anticipated a similar amount today depending on the weather.
  • Rain over the next several days were expected to curtail work until Sunday.